40.25g of glauber's salt is dissolved in water to obtain 500ml of solution of density 1077.5 g dm3 the molality of Na2SO4 in solution is about 

Grade:11

1 Answers

Yash Srivastava
5 years ago
 
Moles of Na2So4.10H20 = 40.25/322 = 0.125 moles 
 
When one mole this is added to water we get one mole Na2So4 and 10 moles of H2o 
So the moles of Na2So4 is 0.125 in the solution
 
Now we find the weight of the solvent
The water getting dissociated later also becomes solvent so 1.25 moles of H20 weighing 22.50 gm
 
Density = 1077.5g/dm3
 
Volume of solution = 0.5 dm3
 
Mass of solution = 538.75g
Mass of solvent = mass of solution – mass of salt added + water dissociated = 521gm
 
Calculating molality
Moles / mass of solvent in kg
0.125/0.521 =0.2399 molal